# FiCo-ITR: bridging fine-grained and coarse-grained image-text retrieval for comparative performance analysis

### v1

@@ 4 Experiment methodology
Datasets. We use MS-COCO and Flickr30K as primary benchmark datasets. MS-COCO contains 123 287 images, each with five human-annotated captions and multi-label annotations across 80 semantic categories. Flickr30K comprises 31K images, also with five human-annotated captions per image, focusing on human-object interactions. To enhance the comparability of the results, the toolkit provided through our FiCo-ITR library implements the Karpathy [50] split uniformly across all evaluated models for consistency. This split allocates 5K images for testing and 5K for validation in MS-COCO, with the remainder used for training. For Flickr30K, 1K images are designated for testing and validation each, with the remaining 29K samples used for training. Due to Flickr30K not containing semantic category labelling, we employ the label generation module of FiCo-ITR enabled by Q2L [48] to generate semantic category labelling for it. These generated labels are available for use within the provided FiCo-ITR repository.
[…]
Evaluation metrics. We use the following configurations of the metrics implemented in the FiCo-ITR library for our experiments: Instance-level retrieval performance is assessed using $R@$1, $R@$5, and $R@$10. Additionally, we explore $R@$50, $R@$100, and $R@$200 for the CG models to examine the viability of their use case as first-screening steps where their top-k results are passed to a FG model for further reranking. For category-level retrieval, we employ mAP@10, mAP@100, and mAP@$N$ (where $N$ is the total number of retrieval candidates), offering a comprehensive view of performance across increasing retrieval depths while having several relevant samples available. To further illustrate the trade-off between precision and recall in category-level retrieval, we include 11-point interpolated precision-recall curves.
@@ 5.1 Instance-level retrieval results
Model comparison and attention mechanisms. The recall performance difference between models employing query-time attention (BLIP-2, X-VLM, IMRAM, SCAN) and those that do not is notable for its limited magnitude. In the case of pretrained models, the state-of-the-art fusion-reranking model BLIP-2 outperforms the dual-encoder model BEIT-3 by a moderate R@1 difference of 6.4 and 6.9 on the MS-COCO (i $\rightarrow$ t) and (t $\rightarrow$ i) tasks, respectively. Similarly, among non-pretrained models, the cross-attention model IMRAM outperforms the global-representation model VSRN by a modest R@1 difference of 4.2 for the MS-COCO (i $\rightarrow$ t) task and 1.6 for (t $\rightarrow$ i) task. Whether these moderate improvements in retrieval performance justify the use of query-time attention is a consideration that the experiments in Section 5.3 aim to inform.
@@ 5.2 Category-level retrieval results
[…]
@@ 5.3 Scaling encoding, storage, and attention
Encoding Time and Storage Cost. Table 5 shows the encoding time and storage requirements for increasing data volumes for the selected models. The experiments were constrained by an 80GiB GPU memory threshold; runs exceeding this limit due to the size of the final or intermediate embeddings did not finish (DNF). For instance, BEIT-3, despite having the most compact final embeddings among VLP models, required 8.12GiB per 1K/5K images/captions to compute intermediate embeddings, causing its 10K/50K encoding run to exceed available GPU memory. These experiments used original model implementations without custom batching to ensure consistency and reflect practical limitations users might encounter with similar hardware. This consideration highlights the storage efficiency of cross-modal hashing models, which require only 0.36GiB to store 1M/5M image/text embeddings. For query-time attention models, Table 5 includes storage costs for original image and text features, which must be retained for the attention step; an important consideration when assessing the practicality of query-time attention in memory-constrained applications. Encoding time generally increased by a factor of 10 from CG to FG methods and again to VLP models, with X-VLM and ADV as exceptions. The rapid encoding times of UCCH and DADH demonstrate the computational efficiency of hash function encoders.
@@
Table 5: Encoding time cost for encoding 1K/5K, 10K/50K, 100K/500K and 1M/5M images/text in seconds and storage costs for holding 1M/5M image/text embeddings and features. DNF: Did not finish due to memory overflow. Features n/a: The model does not require original features during query time
Attention time. Attention in this context refers to cross-attention in the case of IMRAM and SCAN and transformer fusion-based reranking for BLIP-2 and X-VLM, both processes which are computed at query time. Due to the cross-attention mechanism used by IMRAM and SCAN attending to all possible image-text matches, assuming an equal rate of increase in the number of queries and retrieval candidates, the computational complexity of the mechanism approximates $O(n^{2})$. The k value for the reranking shard threshold for BLIP-2 and X-VLM was kept at $k=128$ irrespective of the increase in sample size, which leads to a $O(n)$ increase in attention computation time. However, if the k value were also scaled to the sample size, the computational complexity would approximate $O(n^{2})$. Such attention mechanisms, therefore, may be impractical for large-scale applications where query latency is critical.
@@
Table 6: Time cost for computing cross-attention or transformer self-attention for 1K/5K, 10K/50K and 100K/500K images/text.
@@ 5.4 Scaling similarity search with FAISS
The final step in retrieval, similarity search, is typically independent of the encoding model. Assuming no query-time attention or reranking, the model’s task is complete once samples are encoded. These encoded samples are then typically passed to a separate specialised similarity search implementation. In that case, the only factors affecting similarity search time are the embedding type (real-valued continuous or bitwise hash code embeddings) and the embedding dimensions. To explore the practical implications of using CG binary hash codes compared to FG continuous embeddings, we employ Facebook AI Similarity Search (FAISS) [52], a robust and widely adopted [53, 54, 55] similarity search implementation offering various indexes for both exhaustive search and Approximate Nearest Neighbour Search (ANNS). By using FAISS, we transition from model-specific comparisons to a generalised evaluation of embedding types within an industry-standard framework. The experiments were conducted using the following four indexes:
@@ 6 Key findings and discussion
Attention mechanisms. State-of-the-art benchmark recall results are achieved by models such as BLIP-2, which employ fusion-encoder reranking to refine search results. However, the computational load associated with such query-time attention mechanisms is difficult to justify for practical use. This inefficiency stems from attention mechanisms attending to every possible image-text pair. In scenarios where the increase of images and text samples scale at the same rate, the computational complexity of such attention mechanisms approximates $O(n^{2})$. Therefore, for large-scale retrieval applications where storage concerns can be addressed, dual-encoder VLP models which do not have query-time attention represent the most practical architecture type for instance-level retrieval.
@@ 6.1 Limitations
The scalability experiments were conducted through synthetic duplication of standard benchmark datasets. While this approach offers insights into the computational scalability of the evaluated models, it does not account for potential degradation in retrieval performance as the retrieval candidate pool increases.
Our study primarily focused on existing models and search implementations, aiming for a comprehensive view of the current practical state of the field. This approach aligns with the experiences of practitioners who may rely on publicly available tools and implementations, providing insights into the performance characteristics of widely accessible models and search techniques. However, the application of customised batching implementations within the selected models, as well as custom hardware-level optimisations of CG bitwise operations—both aspects likely to be implemented in production-level use cases—could offer more insightful perspectives on the performance of proprietary production-level pipelines.

### v2

@@ 4 Experiment methodology
Datasets. We use MS-COCO and Flickr30K as primary benchmark datasets. MS-COCO contains 123 287 images, each with five human-annotated captions and native multi-label annotations across 80 semantic categories, making it particularly suitable for evaluating both instance-level and category-level retrieval approaches. MS-COCO’s adoption as a standard benchmark across both FG and CG communities [gan2022vision, luo2023survey] further enables fair cross-methodology comparisons. Flickr30K comprises 31K images, also with five human-annotated captions per image, focusing on human-object interactions. To enhance the comparability of the results, the toolkit provided through our FiCo-ITR library implements the Karpathy [karpathy2017deep] split uniformly across all evaluated models for consistency. This split allocates 5K images for testing and 5K for validation in MS-COCO, with the remainder used for training. For Flickr30K, 1K images are designated for testing and validation each, with the remaining 29K samples used for training. Due to Flickr30K not containing semantic category labelling, we employ the label generation module of FiCo-ITR enabled by Q2L [liu2021query2label] to generate semantic category labelling for it. These generated labels are available for use within the provided FiCo-ITR repository.
Model selection. The model selection process considered both academic impact and architectural diversity while ensuring practical reproducibility. From recent surveys and benchmarks [luo2023survey, gan2022vision], we identified representative models across the key architectural approaches in ITR (as detailed in Table 1). Our model selection thus aims to span the architectural spectrum: For VLP models, we include the fusion-encoder ViLT [kim2021vilt], dual-encoder BEiT-3 [wang2023image], and hybrid dual+fusion rerank models BLIP-2 [li2023blip] and X-VLM [zeng2022multi]. Traditional FG models are represented through both local+global attention (IMRAM [chen2020imram], SCAN [lee2018stacked]) and global-only attention (VSRN [li2019visual]) approaches, while CG models cover the main paradigms: supervised (DADH [bai2020deep]), unsupervised (UCCH [hu2022unsupervised]), and quantisation-based (ADV [li2021efficient]) approaches. Some efficiency-focused VLP models of interest (LightningDot [sun2021lightningdot], VLDeformer [zhang2022vldeformer], and HiVLP [chen2022hivlp]) could not be included due to practical constraints such as inaccessible data dependencies or lack of public implementations. To partially address this gap, we implement dual-encoder-only variants of BLIP-2 and X-VLM to promote efficiency ($\text{BLIP-2}_{\text{NF}}$ and $\text{X-VLM}_{\text{NF}}$). Additionally, we evaluate ADV in two configurations: as a CG model with 64-bit hash codes and as a FG model with 2048-bit codes. Table 1 provides a detailed overview of the architectural characteristics and computational requirements across our model selection.
Evaluation metrics. We use the following configurations of the metrics implemented in the FiCo-ITR library for our experiments: Instance-level retrieval performance is assessed using R@R@1, R@R@5, and R@R@10. Additionally, we explore R@R@50, R@R@100, and R@R@200 for the CG models to examine the viability of their use case as first-screening steps where their top-k results are passed to a FG model for further reranking. For category-level retrieval, we employ mAP@10, mAP@100, and mAP@$N$ (where $N$ is the total number of retrieval candidates), offering a comprehensive view of performance across increasing retrieval depths while having several relevant samples available. To further illustrate the trade-off between precision and recall in category-level retrieval, we include 11-point interpolated precision-recall curves.
@@ 5.1 Instance-level retrieval results
Model comparison and attention mechanisms. The recall performance difference between models employing query-time attention (BLIP-2, X-VLM, ViLT, IMRAM, SCAN) and those that do not is notable for its limited magnitude. In the case of pretrained models, the state-of-the-art fusion-reranking model BLIP-2 outperforms the dual-encoder model BEIT-3 by a moderate R@1 difference of 6.4 and 6.9 on the MS-COCO (i $\rightarrow$ t) and (t $\rightarrow$ i) tasks, respectively. The effectiveness of dual-encoder architectures is further evidenced by BEIT-3 achieving superior performance across all evaluation metrics compared to the full fusion-based ViLT, demonstrating that competitive retrieval performance can be attained without extensive inter-modal interaction at query time. Similarly, among non-pretrained models, the cross-attention model IMRAM outperforms the global-representation model VSRN by a modest R@1 difference of 4.2 for the MS-COCO (i $\rightarrow$ t) task and 1.6 for (t $\rightarrow$ i) task. Whether these moderate improvements in retrieval performance justify the use of query-time attention is a consideration that the experiments in Section 5.3 aim to inform.
@@ 5.2 Category-level retrieval results
mAP@N challenge and model size. In contrast to the relative ease of scoring well on the mAP@10 metric, the mAP@N metric presents a more substantial challenge. Unlike instance-level retrieval, performance on this metric does not appear to align with the size of the model being used. For example, on the MS-COCO dataset, the state-of-the-art FG VLP $\text{BLIP-2}_{\text{NF}}$ model achieves the second lowest category-level mAP@N score of 47.9% on the (i $\rightarrow$ t) task, while the coarse model UCCH attains the highest score of 60.7% (Table 4). This suggests that FG low-level semantics may introduce noise when high-level category semantics is all that is needed for the task.
@@
Fusion-encoder global similarity limitation. The fusion-encoder approach implemented by ViLT is less suitable for category-level retrieval due to both its architectural design and training approach. ViLT is trained using a binary matching approach where the model learns to distinguish correct pairings from a limited set of negative samples [kim2021vilt]. This setup trains the model to give high scores only to exact correct pairing while treating all other combinations as non-matches. However, for category-level retrieval, we need a model that can identify many relevant items that share semantic similarities with the query, not just find the single exact match. As a result, fusion-based models like ViLT, while effective at specific instance-level retrieval, perform poorly at ranking more generally similar items at the category level.
@@ 5.3 Scaling encoding, storage, and attention
Encoding time and storage cost. Table 5 shows the encoding time and storage requirements for increasing data volumes for the selected models. The experiments were constrained by an 80GiB GPU memory threshold; runs exceeding this limit due to the size of the final or intermediate embeddings did not finish (DNF). For instance, BEIT-3, despite having the most compact final embeddings among VLP models, required 8.12GiB per 1K/5K images/captions to compute intermediate embeddings, causing its 10K/50K encoding run to exceed available GPU memory. These experiments used original model implementations without custom batching to ensure consistency and reflect practical limitations users might encounter with similar hardware. This consideration highlights the storage efficiency of cross-modal hashing models, which require only 0.36GiB to store 1M/5M image/text embeddings. For query-time attention models, Table 5 includes storage costs for original image and text features, which must be retained for the attention step; an important consideration when assessing the practicality of query-time attention in memory-constrained applications. Encoding time generally increased by a factor of 10 from CG to FG methods and again to VLP models, with X-VLM and ADV as exceptions. The rapid encoding times of UCCH and DADH demonstrate the computational efficiency of hash function encoders. ViLT’s end-to-end fusion architecture computes similarity scores directly without encoding intermediate embeddings, requiring full model inference for each new query and preventing the possibility of offline embedding storage and reuse. Consequently, ViLT’s computational costs are not represented in Table 5, which focuses on one-time encoding costs for offline storage; instead, its query-time costs are analyzed in the following segment on attention mechanisms.
@@
Table 5: Encoding time cost for encoding 1K/5K, 10K/50K, 100K/500K and 1M/5M images/text in seconds and storage costs for holding 1M/5M image/text embeddings and features. ViLT’s encoding time and embedding storage are not applicable due to its end-to-end fusion architecture, which computes similarity scores directly without encoding embeddings. For ViLT’s computational costs, see Table 6. DNF: Did not finish due to memory overflow. Features n/a: The model does not require original features during query time
Attention time. Attention in this context refers to end-to-end fusion for ViLT, fusion-based reranking for BLIP-2 and X-VLM and cross-attention for IMRAM and SCAN; processes which are computed at query time. Due to the end-to-end fusion and cross-attention mechanisms used by ViLT, IMRAM and SCAN attending to all possible image-text pairs, the computational complexity of the mechanism is $O(n\times m)$, where n is the number of queries and m the number of retrieval candidates, quickly making retrieval prohibitively expensive computationally. This computational burden has driven the field towards dual-encoder architectures, which avoid query-time attention by computing and storing embeddings independently for each modality. For fusion-reranking, the time complexity is $O(n\times k)$, where k is the reranking shard threshold, which leads to improved costs relative to full attention. In the case of BLIP-2 and X-VLM, we kept $k=128$ irrespective of the increase in sample size, which leads to a linear increase in attention computation time. However, if the k value were also scaled to the number of retrieval candidates, the computational complexity would approximate $O(n\times m)$. Such attention mechanisms, therefore, may be impractical for large-scale applications where query latency is critical.
@@
Table 6: Time cost for computing end-to-end fusion, fusion re-ranking, cross-attention 1K/5K, 10K/50K and 100K/500K images/text.
@@ 5.4 Scaling similarity search with FAISS
The final step in retrieval, similarity search, is typically independent of the encoding model. Assuming no query-time attention or reranking, the model’s task is complete once samples are encoded. These encoded samples are then typically passed to a separate specialised similarity search implementation. In that case, the only factors affecting similarity search time are the embedding type (real-valued continuous or bitwise hash code embeddings) and the embedding dimensions. To explore the practical implications of using CG binary hash codes compared to FG continuous embeddings, we employ Facebook AI Similarity Search (FAISS) [johnson21billion], a robust and widely adopted [pizzi2022self, barrault2023seamlessm4t, thakur2021beir] similarity search implementation offering various indexes for both exhaustive search and Approximate Nearest Neighbour Search (ANNS). By using FAISS, we transition from model-specific comparisons to a generalised evaluation of embedding types within an industry-standard framework. The experiments were conducted using the following four indexes:
@@ 6 Key findings and discussion
Attention mechanisms. State-of-the-art benchmark recall results are achieved by models such as BLIP-2, which employ fusion-encoder reranking to refine search results. However, the computational load associated with such query-time attention mechanisms is difficult to justify for practical use in retrieval applications. This inefficiency stems from fusion encoders attending to every possible image-text pair with a computational complexity of $O(n\times m)$, where n is the number of queries and m is the number of retrieval candidates. Unlike retrieval tasks where fusion encoders must process all query-candidate pairs, in applications like VQA or image captioning, the fusion encoder only needs to attend to a single image-text pair at a time (e.g. the given question and image), resulting in $O(n)$ complexity. This makes fusion encoders particularly well-suited for these single-pair reasoning tasks[ishmam2024image, stefanini2022show], but prohibitively expensive for large-scale retrieval. Therefore, for retrieval applications where storage concerns can be addressed, dual-encoder VLP models which do not have query-time attention represent the most practical architecture type for instance-level retrieval.
@@ 6.1 Limitations
Our proposed FiCo-ITR library and experiments identify new insights into the tradeoffs of FG and CG ITR models, however, we acknowledge certain limitations. CG models are commonly benchmarked on datasets which contain hashtag collections as text samples (e.g. MIR-Flickr25K [huiskes2008mir] or NUS-WIDE [chua2009nus]). In contrast, FG models require full sentences to enable instance-level retrieval. There are currently no benchmark ITR datasets that contain both full sentences and hashtag collections as their text samples. This limitation presents an opportunity for an extension of our work, where an image captioning model could be used to generate full sentences for MIR-Flickr25K or NUS-WIDE to enable further evaluations.
Our scalability experiments quantified the computational costs associated with increasingly large retrieval sets. To achieve this, we used data duplication to grow our datasets. However, the potential degradation of retrieval performance as the size of the retrieval set increases could not be measured through this approach. Large-scale ITR benchmarks which would enable this investigation do not exist in the field. Currently available large-scale image-text datasets (e.g. LAION-5B [schuhmann2022laion], CC13 [changpinyo2021conceptual]) are built for vision-language pretraining and have data imbalance and noise, which makes them unsuitable for benchmarking purposes. Our future work will target creating large-scale ITR benchmark datasets which maintain data quality and balance.
Our study could be extended in several directions: Our experiments focused on a cross-methodology comparison of FG and CG approaches. However, further within-methodology experiments could inform future work, such as implementing FG model ensembles or identifying the most suitable CG architecture for real-world use cases. Additional experiments could examine the generalisability of our findings across different data distributions, application domains, and model sensitivity to different types of data. Further efficiency gains could be achieved through hardware optimisations and custom batching solutions.